The pressure of a gas at constant volume is 90 mm Hg at 20 degree celcius, calculate its pressure at 70 degree celcius to the nearest whole number
1
Expert's answer
2020-07-02T17:13:35-0400
"\frac{p_1}{T_1}=\frac{p_2}{T_2}\\\frac{90}{273+20}=\frac{p_2}{273+70}\\p_2=105\text{ mm Hg}"